OpenStudy (anonymous):

simplify the expression: t^2-81 / t-9

OpenStudy (anonymous):

(t+9)(t-9)/(t-9) =(t+9)

OpenStudy (anonymous):

i know that the bottom will simplify to 1(t-9) ut i'm not sure about the numerator but i know it willhave (t-9)

OpenStudy (anonymous):

ohhh ok i understand!

OpenStudy (anonymous):

in the numerator use the identity (x^2-a^2)=(x+a)(x-a) and 81=9^2